# On the number of permutation-twisted dot products

### v1

@@ Abstract.
For distinct real numbers $a_{1},\ldots,a_{n}$ and distinct real numbers $b_{1},\ldots,b_{n}$, consider the sum $S=\sum_{i=1}^{n}a_{i}b_{\pi(i)}$ as $\pi$ ranges over the permutations of $[n]$. We show that this sum always assumes at least $\Omega(n^{3})$ distinct values, which is optimal. This “support” bound complements recent work of Do, Nguyen, Phan, Tran, and Vu on the anticoncentration properties of $S$ when $\pi$ is chosen uniformly at random.
@@ 1. Introduction
which is the “dot product of $\mathbf{a},\mathbf{b}$ twisted by $\pi$”. Motivated by Littlewood–Offord theory and a geometric problem of Pawlowski [P24], Do, Nguyen, Phan, Tran, and Vu [DNPTV25] recently raised the problem of studying the anticoncentration properties of the random variable $S(\mathbf{a},\mathbf{b};\pi)$ where $\pi$ is chosen uniformly at random. One of their main results states that if the $a_{i}$’s are all distinct and the $b_{i}$’s are all distinct, then each possible value of the sum is assumed with probability at most $O(n^{-5/2}\log n)$; this bound is optimal up to the logarithm (which likely can be removed). This work has prompted a small flurry of activity [ABDV26, HPZ26].
@@
of all possible sums as $\pi$ ranges; this is of course precisely the support of Do, Nguyen, Phan, Tran, and Vu’s random variable $S(\mathbf{a},\mathbf{b};\pi)$. Our main result is that if the $a_{i}$’s are all distinct and the $b_{i}$’s are all distinct, then there are at least $\Omega(n^{3})$ distinct sums.
@@
Our proof of Theorem 1.1 is elementary and fairly short. As it turns out, we are able to obtain $\Omega(n^{3})$ distinct sums using a somewhat small set of permutations, namely, the permutations obtained as an involution times a fixed “base” permutation. Pohoata [P26] has recently obtained an independent proof of Theorem 1.1 using more complicated methods.
@@ 2.1. Constructing the involutions
The first step is identifying a family of disjoint pairs $(j_{i},k_{i})$ of elements of $[n]$ such that the corresponding differences $a_{k_{i}}-a_{j_{i}}$ “see” the various scales of $\mathbf{a}$; and likewise for $\mathbf{b}$. We achieve this iteratively, as follows.
@@
Write $i_{1}<i_{2}<\cdots<i_{u+v-1}$ for the elements of $\mathcal{U}_{u+v}$ between $j_{u+v}$ and $k_{u+v}$. All of these elements are also contained in $\mathcal{U}_{u}$ and $\mathcal{U}_{v}$, so the greedy nature of our inductive procedure guarantees that $x(u)\leq a_{i_{u}}-a_{j_{u+v}}$ and $x(v)\leq a_{k_{u+v}}-a_{i_{u}}$, whence $x(u)+x(v)\leq a_{k_{u+v}}-a_{j_{u+v}}=x(u+v)$. ∎

### v2

@@ Abstract.
For distinct real numbers $a_{1},\ldots,a_{n}$ and distinct real numbers $b_{1},\ldots,b_{n}$, consider the sum $S=\sum_{i=1}^{n}a_{i}b_{\pi(i)}$ as $\pi$ ranges over the permutations of $[n]$. We show that this sum always assumes at least $\Omega(n^{3})$ distinct values, which is optimal. This “support” bound complements recent work of Do, Nguyen, Phan, Tran, and Vu, and of Hunter, Pohoata, and Zhu on the anticoncentration properties of $S$ when $\pi$ is chosen uniformly at random.
@@ 1. Introduction
which is the “dot product of $\mathbf{a},\mathbf{b}$ twisted by $\pi$”. Motivated by Littlewood–Offord theory and a geometric problem of Pawlowski [P24], Do, Nguyen, Phan, Tran, and Vu [DNPTV25], and independently Hunter, Pohoata, and Zhu [HPZ26] recently raised the problem of studying the anticoncentration properties of the random variable $S(\mathbf{a},\mathbf{b};\pi)$ where $\pi$ is chosen uniformly at random. One of their main results states that if the $a_{i}$’s are all distinct and the $b_{i}$’s are all distinct, then each possible value of the sum is assumed with probability at most $O(n^{-5/2}\log n)$; this bound is optimal up to the logarithm (which likely can be removed). The paper [ABDV26] has rounded out this small flurry of recent activity.
@@
of all possible sums as $\pi$ ranges; this is of course precisely the support of the random variable $S(\mathbf{a},\mathbf{b};\pi)$. Our main result is that if the $a_{i}$’s are all distinct and the $b_{i}$’s are all distinct, then there are at least $\Omega(n^{3})$ distinct sums.
@@
Our proof of Theorem 1.1 is elementary and fairly short. As it turns out, we are able to obtain $\Omega(n^{3})$ distinct sums using a somewhat small set of permutations, namely, the permutations obtained as involutions multiplied by a fixed “base” permutation. Pohoata [P26] has recently obtained an independent proof of Theorem 1.1 using more complicated methods.
@@ 2.1. Constructing the involutions
The first step is identifying a family of disjoint pairs $(j_{i},k_{i})$ of elements of $[n]$ such that the corresponding differences $a_{k_{i}}-a_{j_{i}}$ “see” the various scales of $\mathbf{a}$, and likewise for $\mathbf{b}$. We achieve this iteratively, as follows.
@@
Write $i_{1}<i_{2}<\cdots<i_{u+v-1}$ for the elements of $\mathcal{U}_{u+v}$ between $j_{u+v}$ and $k_{u+v}$. All of these elements are also in $\mathcal{U}_{u-1}$ and $\mathcal{U}_{v-1}$, so the greedy nature of our inductive procedure guarantees that $x(u)\leq a_{i_{u}}-a_{j_{u+v}}$ and $x(v)\leq a_{k_{u+v}}-a_{i_{u}}$, whence $x(u)+x(v)\leq a_{k_{u+v}}-a_{j_{u+v}}=x(u+v)$. ∎